Electronics > Repair

HP 3458A - Fails Slave Test - Convergence

<< < (4/16) > >>

Kleinstein:
The convergance failure / salve test fail and hang up state point a little towards the ADC board (A3).  Here known failing parts are the already mentioned comparators, the fiber link (e.g. loose connector) and also the worst case U180 failure.

A drifting signal in the voltage reading mode cold be something like a porblem on A1 with the input path (e.g. relays or as TiN just had it with the protection resistors).  A open path there can make the amplifier slowly dirft. As the readings in current and AC mode seem stable, this points in this direction. Of cause just an open input (e.g. switched to rear) could also cause drift and an overlaod reading with ohm tests.

The odd trigger behavior could be as simple as a odd configuration still remembered in the buffered RAM.

Ade911:
Kleinstein thank you for your feedback and advice. I want to believe that the A3 Bd is OK only because  it does measure a value  plus its expensive to replace.

Plan of action,

Purchase new memory devices and try to copy over the existing data.
Break the loop on the A1 Bd and see where the rise in voltage is coming from around Q106, Q107, Q108.

I read that the original FETS are obsolete, has anyone found an alternative device just in case one is leaking ?

Many tx Ade.

Ade911:
Sorry meant to ask, as the Dallas models are now obsolete which version should i purchase.

Also what type of reader , programmer is recommended wih out spending mega bucks.

Many tx Ade.

Ade911:
Update

Purchased a GQ-4x4 programmer , removed all Dallas devices and copied data to new ones , replacing in the meter , now suffer from needing all calibrations , not sure if the copy was OK although the programmer said they were - really confused over this. BTW the programmed sometime [if you programmed twice or  ore] would randomly say there was a problem with the Dallas chips , not sure if others have had similar problems with these programmers.

In parallel I found the output of U105 which drives analog switches on A1 was floating all over the place , what ever i tried on the input had no effect on the output so removed , need to wait for a new device but for now running LM709 which is not ideal but I dont mind as the out put which drives U104 is grounded anyhow so I could actually run without the device I believe at present.

As suggested I also checked all the fibre connections which seem to work OK well I believe they do , scoping one end and the other the signals jump up and down at the same time.

So now things get worse, the system was trying to calibrate DCV before I checked the fibre links, it was sampling but now nothing.

It only samples when I activate TARM and in KOHMS although it shows overload.

TP100 is also slowly drifting up to the supply rail voltage, Ive replaced U103, U108 and still no joy.

The input to Q103 is stable , zero volts if i recall , if I remove U103 everything is fine and I don't see any drift so could it be the FET Q107 Q108 well not sure. As I don't have spare and Im not sure if we can get replacements - SOMEONE PLEASE CONFIRM IF IM WRONG as Im scared to remove them.

Can someone please also confirm where the calibration 10V comes from as well , again I'm confused to know where its generated and fed from , you normally see in the screen when sampling but I don't.

Starting to wonder if Ive started on something i cant repair.


Kleinstein:
There are too many, more random changes of parts and to little measurements / checks before. This way lots of parts get touched and possible new problems added.

The part around U105 and U104 should not effect the normal voltage readings in the 10 V or 1 V ranges, even of the part is bad or possibly missing. So no need to worry about them for the start.
These 2 may effect the 100 mV range and especially the 4 wire low value resistance ranges.

Removing U103 (LM399) would leave more than 1 FET switch active at the same time for the gain setting. Worst case this may damage things, though chances are the parts would survive.

For debugging the input amplifier it would help to start with the non AZ mode, though with a stable 0 V at the input it would not make a difference.
For the 10 V range pins 1 and 2 of U103 should be at the negative rail, some -18 V.
Pin 14 should follow TP100 = the amplifier output.
TP101 as the buffered feedback signal is also of interest.

The data for the 10 V calibration constant should come from the CAL ram chip.
A 10 V signal is used in the ACAL procedure in the calibration of the HV divider. Where that signal is comming from is a good question. I don't know for sure, but it could come via the shunt chain and Ohms source.

Navigation

[0] Message Index

[#] Next page

[*] Previous page

There was an error while thanking
Thanking...
Go to full version
Powered by SMFPacks Advanced Attachments Uploader Mod